# R-MTLLMF: Resilient Multi-Task Large Language Model Fusion at the Wireless Edge

### v2

@@ I Introduction and Motivation
Future 6G wireless networks will enable reliable, low-latency, and intelligent communications to support critical technologies such as artificial intelligence (AI) [1], distributed and collaborative machine learning (ML) [2], and intelligent edge computing [3]. Recent advancements in large language models (LLMs) in particular have motivated several use cases for the wireless industry, including network design, resource allocation, and standardization [4]. In this generative AI vision of the future, distributed agents will play a central role in fine-tuning, sharing and merging LLMs at the wireless edge [5]. For example, [6] presents a framework where distributed agents use LLMs for reasoning and collaboration in wireless networks, supporting knowledge-based edge device queries.
@@ III R-MTLLMF: AI-Driven Resilience Against Adversarial Noise
  • Task Vector Aggregation with Frozen LLM Parameters: Given the severity of noise at inference time, we freeze sensitive position, patch and class embeddings of the base model $\theta_{\text{base}}$, and reload them into $\theta_{\text{MTLLM}}$ after task vector aggregation. Freezing these parameters ensures that fundamental data representations remain intact, reducing the risk of noise-induced shifts in essential encodings. This is a valid approach since fine-tuning pre-trained models mostly affects the self-attention layers, and since $\theta_{\text{base}}$ remains identical across tasks [5]. Thus, by preserving embeddings, we maintain the integrity of structural representations learned during pre-training, thereby stabilizing the model’s output.
  • Few-Shot Realignment: To correct for noise-induced perturbations in self-attention parameters, we apply a few-shot fine-tuning step to realign the MTLLM on a small, representative data subset (typically 10 samples per class). This can be sourced from episodic memory, as in [13], or a public dataset for privacy considerations, and can be performed at the MTMF BS. By realigning the attention mechanism on few samples, the model restores the contextual relationships between tokens that may be distorted due to noise. Drawing on insights from task vector research in [8, 5], this heuristic relies on the notion that noise, if not substantial, leaves the underlying task vector directionality of $\hat{\tau}_{q}$ largely intact, allowing the model’s performance to be recovered with minimal fine-tuning. Thus, realignment addresses moderate perturbations, restoring the model’s reasoning capabilities.
@@ IV Simulation Results and Analysis
We follow the experimental setup in in [9] and fine-tune OpenAI’s CLIP ViT-B/16 ViT-LLM [14] for 2000 iterations with batch size 128 on eight different image classification datasets (tasks): Cars [15], DTD [16], EuroSAT [17], GTSRB [18], MNIST [19], RESISC45 [20], SUN397 [21], and SVHN [22]. We then compute the task vectors and construct the MTLLM with parameters $\theta_{\text{MTLLM}}$ as defined in (6) with heuristic values for $\lambda_{N}$ from [8]. We vary the number of tasks between 2 and 8, and, for each case, we generate all possible task combinations and evaluate the MTMF cross-task performance by measuring accuracy on each dataset, normalized to the individual fine-tuning performance. We report the average of the normalized accuracies as in [9]. To simulate ideal and adversarial channel noise conditions for sum rate (SR) and strongest user (SU) per (P1) and (P2), we consider the MIMO setup from Section II with $|\mathcal{Q}|=8$ single-antenna users, each of which fine-tunes on one of above datasets, and $N_{R}=16$ BS receive antennas. We set the maximum transmit power to 0.1 mW. We initialize the user positions randomly between 100 to 1000 meters from the BS and employ 64-QAM modulation. To induce parameter deviation from post-decoding errors $\tilde{e}_{q}$, we compute the MSE and add Gaussian noise with variance proportional to it to the task vectors, resulting in $\hat{\tau}_{q}=\tau_{q}+\tilde{\epsilon}_{q}$. We apply R-MTLLMF with few-shot realignment, using 10 random samples per class as in [13].
@@ IV-1 R-MTLLMF Results
Results for MTMF with 8 tasks are shown in Figure 2 where shaded areas represent the variance in accuracies for all task combinations. R-MTLLMF achieves close-to-baseline performance for the ideal transmission case, having 89.1% normalized average accuracy compared to the 91.2% baseline when aggregating all 8 tasks. In contrast, R-MTLLMF has equally degraded performance for both SR and SU worst-case scenarios with 54.8% and 53.3% accuracies, respectively. This is due to the significantly increased noise during task aggregation. Whereas the ideal scenario results in a signal-to-noise-ratio (SNR) of around $-16.2$ dB with an average MSE of $0.013$ across all 8 users, results for SU (SR) yield an SNR of -21.7 dB (-27.3 dB) with an average MSE of $0.16$ ($0.21$). The severe increase in noise shows some limitations of our AI-based framework where neither parameter freezing nor few-shot realignment can fully restore the model’s classification capabilities as the cross-task interference is too high. However, even in this worst-case, R-MTLLMF ensures significantly better model performance (9-14 times better) as compared to scenarios without protection. More concretely, unprotected SR and SU scenarios lead to catastrophic perturbances with average normalized accuracies between 4% and 6% across all tasks, resulting in worst-case single- and multi-task performance, which ultimately renders MTMF ineffective. Surprisingly, even the unprotected ideal case results in subpar performance with total aggregation accuracy of 36.3% due to noise at inference time. Furthermore, accuracies seem to be generally non-increasing with the number of tasks, except for R-MTLLMF in the ideal case. This indicates that the perturbed model is unable to generalize across tasks. In addition, the heuristic choice for $\lambda_{N}$ may be suboptimal, indicated by random improvements and deteriorations. This signifies the importance of resilient aggregation strategies in distributed inference scenarios, where noise, even if small, may be detrimental.
@@ IV-3 Ablation Studies
In summary, MTMF improves the average performance as the number of tasks increases without the need for constant re-training. Furthermore, our R-MTLLMF framework provides an effective AI-based MTMF resilience framework for most
@@
typical wireless noise scenarios. It leads to increased task vector orthogonality and thus improves multi-task performance, achieving close-to-baseline results. However, its effectiveness diminishes under extreme adversarial noise. This suggests that AI-based resilience strategies alone may be insufficient for worst-case attacks, and we recommend augmenting R-MTLLMF with complementary physical layer techniques. Despite this limitation, R-MTLLMF still outperforms unprotected setups by a factor of 9-14 under worst-case noise, making it a substantial improvement over standard MTMF.

### v3

@@ I Introduction and Motivation
Future 6G wireless networks will enable reliable, low-latency, and intelligent communications to support critical technologies such as artificial intelligence (AI) [1], distributed and collaborative machine learning (ML) [2], and intelligent edge computing [3]. Recent advancements in large language models (LLMs) in particular have motivated several use cases for the wireless industry, including network design, resource allocation, and standardization [4]. In this generative AI vision of the future, distributed agents will play a central role in fine-tuning, sharing and merging LLMs at the wireless edge. For example, [5] presents a framework where distributed agents use LLMs for reasoning and collaboration in wireless networks, supporting knowledge-based edge device queries.
@@ III R-MTLLMF: AI-Driven Resilience Against Adversarial Noise
  • Task Vector Aggregation with Frozen LLM Parameters: Given the severity of noise at inference time, we freeze sensitive position, patch and class embeddings of the base model $\theta_{\text{base}}$, and reload them into $\theta_{\text{MTLLM}}$ after task vector aggregation. Freezing these parameters ensures that fundamental data representations remain intact, reducing the risk of noise-induced shifts in essential encodings. This is a valid approach since fine-tuning pre-trained models mostly affects the self-attention layers, and since $\theta_{\text{base}}$ remains identical across tasks [7, 8]. Thus, by preserving embeddings, we maintain the integrity of structural representations learned during pre-training, thereby stabilizing the model’s output.
  • Few-Shot Realignment: To correct for noise-induced perturbations in self-attention parameters, we apply a few-shot fine-tuning step to realign the MTLLM on a small, representative data subset (typically 10 samples per class). This can be sourced from episodic memory, as in [12], or a public dataset for privacy considerations, and can be performed at the MTMF BS. By realigning the attention mechanism on few samples, the model restores the contextual relationships between tokens that may be distorted due to noise. Drawing on insights from task vector research in [7], this heuristic relies on the notion that noise, if not substantial, leaves the underlying task vector directionality of $\hat{\tau}_{q}$ largely intact, allowing the model’s performance to be recovered with minimal fine-tuning. Thus, realignment addresses moderate perturbations, restoring the model’s reasoning capabilities.
@@ IV Simulation Results and Analysis
We follow the experimental setup in [8] and fine-tune OpenAI’s CLIP ViT-B/16 ViT-LLM [13] for 2000 iterations with batch size 128 on eight different image classification datasets (tasks): MNIST, Cars [14], DTD [15], EuroSAT [16], GTSRB [17], RESISC45 [18], SUN397 [19], and SVHN [20]. We then compute the task vectors and construct the MTLLM with parameters $\theta_{\text{MTLLM}}$ as defined in (6) with heuristic values for $\lambda_{N}$ from [7]. We vary the number of tasks between 2 and 8, and, for each case, we generate all possible task combinations and evaluate the MTMF cross-task performance by measuring accuracy on each dataset, normalized to the individual fine-tuning performance. We report the average of the normalized accuracies as in [8]. To simulate ideal and adversarial channel noise conditions for sum rate (SR) and strongest user (SU) per (P1) and (P2), we consider the MIMO setup from Section II with $|\mathcal{Q}|=8$ single-antenna users, each of which fine-tunes on one of above datasets, and $N_{R}=16$ BS receive antennas. We set the maximum transmit power to 0.1 mW. We initialize the user positions randomly between 100 to 1000 meters from the BS and employ 64-QAM modulation. To induce parameter deviation from post-decoding errors $\tilde{e}_{q}$, we compute the MSE and add Gaussian noise with variance proportional to it to the task vectors, resulting in $\hat{\tau}_{q}=\tau_{q}+\tilde{\epsilon}_{q}$. We apply R-MTLLMF with few-shot realignment, using 10 random samples per class.
@@ IV-1 R-MTLLMF Results
Figure 2 shows results for 8 tasks where shaded areas represent the variance in accuracies for all task combinations. R-MTLLMF achieves close-to-baseline performance for ideal transmission, having 89.1% normalized average accuracy compared to the 91.2% baseline when aggregating all 8 tasks. In contrast, R-MTLLMF has equally degraded performance for both SR and SU worst-case scenarios with 54.8% and 53.3% accuracies, respectively, due to the significantly increased noise during task aggregation. Whereas the ideal scenario results in a signal-to-noise-ratio (SNR) of $-16.2$ dB with an average MSE of $0.013$ across all 8 users, results for SU (SR) yield an SNR of -21.7 dB (-27.3 dB) with an average MSE of $0.16$ ($0.21$). The severe increase in noise shows some limitations of our AI-based framework where neither parameter freezing nor few-shot realignment can fully restore the model’s classification capabilities as the cross-task interference is too high. However, even in this worst-case, R-MTLLMF ensures significantly better model performance (9-14 times better) as compared to scenarios without protection. More concretely, unprotected SR and SU scenarios lead to catastrophic perturbances with accuracies between 4% and 6%,
resulting in worst-case single- and multi-task performance, which ultimately renders MTMF ineffective. Surprisingly, even the unprotected ideal case results in subpar performance with total aggregation accuracy of 36.3% due to noise at inference time. Furthermore, accuracies seem to be generally non-increasing with the number of tasks, except for R-MTLLMF in the ideal case. This indicates that the perturbed model is unable to generalize across tasks. In addition, the heuristic choice for $\lambda_{N}$ may be suboptimal, indicated by random improvements and deteriorations. This signifies the importance of resilient aggregation strategies in distributed inference scenarios, where noise, even if small, may be detrimental.
@@ IV-3 Ablation Studies
In summary, MTMF improves the average performance as the number of tasks increases without the need for constant re-training. Our R-MTLLMF framework further provides an effective AI-based MTMF resilience framework for most typical wireless noise scenarios. It leads to increased task vector orthogonality and improves multi-task performance, achieving close-to-baseline results. However, its effectiveness diminishes
@@
under extreme adversarial noise. This suggests that AI-based resilience strategies alone may be insufficient for worst-case attacks, including sophisticated jamming that cannot be modeled as such and may lead to catastrophic denial of service [21, 22]. Thus, we recommend augmenting R-MTLLMF with complementary physical layer techniques, potentially enabled by future integrated sensing and communication (ISAC) [9]. Despite this limitation, R-MTLLMF still outperforms unprotected setups by a factor of 9-14 under worst-case noise, making it a substantial improvement over standard MTMF.